Using transformation - solve each equation:
4/9x = 8
n/2.3 = 5.6
3/5n + 15 = 21
13 - 1/3n = 9
3/4n + 22 = 55

Answer by stanbon(48545)

(4/9)x = 8
Multiply both sides by 9/4 to get:
x=72/4=18
----------------
n/2.3 = 5.6
Multiply both sides by 2.3 to get:
n=12.88
--------------
(3/5)n + 15 = 21
(3/5)n=6
Multiply both sides by 5/3 to get:
n=10
--------------
13 - (1/3)n = 9
(1/3)n=3
n=9
------------
(3/4)n + 22 = 55
(3/4)n=33
n=44
Cheers,
Stan H.